KT Televisions, a US firm, has a Chinese subsidiary that manufactures and sells TVs in China.

a. Main input is priced in USD (USD110/unit)

b. All other costs are in RMB (Fixed cost=RMB4M, Variable cost= RMB460/unit).

c. Depreciation = RMB1.3M

d. S0 = RMB6.26/USD

e. Expects to sell 7,000 TVs this year at RMB2,000 each.

f. Tax rate=30%; assuming tax credits are available for immediate use if losses occur

You are required to show all your workings: a) What are the operating cash flows in RMB and dollars?

b) How many unit KT Televisions needs to sell to break-even in operating cash flows in dollars? (Hint: round your answer up to the nearest unit)

Ignore part (a) and part (b) for the following questions and assume that the normal sales are 8,000 Units for the rest of the questions. c) What are your operating cash flows in dollars now?

d) If the spot rate reduces to RMB4.6/USD, KT would like to pass all benefits to his client by reducing selling price. What would be the new selling price that would maintain his profit (operating cash flows in dollars) in Part (c) and would pass all benefits to his client at the same time?
Answer #1
Operating Cash Flow(RMB) RMB
A Number of units sold 7000
B Sales Price per unit 2000
C=A*B Total Sales Revenue          14,000,000
D Input Price per unit=110*6.26 688.60
E=A*D Total Input cost          (4,820,200)
F Unit Variable Cost 460
G=A*F Total Variable Costs          (3,220,000)
H Fixed costs          (4,000,000)
I Depreciation          (1,300,000)
J=C+E+G+H+I Before tax operating income                659,800
K=J*30% Tax expenses              (197,940)
L=J+K After tax Operating Income                461,860
M Add:Depreciation (non cash expense)            1,300,000
N=L+M Operating Cash Flow            1,761,860
Operating Cash Flow(Dollars) Dollar
A Number of units sold 7000
B Sales Price per unit $319.49 (2000/6.26)
C=A*B Total Sales Revenue $2,236,422
D Input Price per unit $110.00
E=A*D Total Input cost -$770,000
F Unit Variable Cost $73.48 (460/6.26)
G=A*F Total Variable Costs -$514,377
H Fixed costs -$638,978 (4million/6.26)
I Depreciation -$207,668 (1.3 million/6.26)
J=C+E+G+H+I Before tax operating income $105,399
K=J*30% Tax expenses -$31,620
L=J+K After tax Operating Income $73,780
M Add:Depreciation (non cash expense) $207,668
N=L+M Operating Cash Flow $281,447
b) Break Even units (Cash flow in dollars)
CM=B-D-F Contribution Margin Per Unit $136.01
FC=-H-I Total Fixed Costs $846,645
Break Even units =Fixed Costs/Unit Contribution margin
BEP=FC/CM Accounting Break Even Point in units 6225
Operating Cash Flow(Dollar) Dollar
A Number of units sold 4044
B Sales Price per unit $319.49
C=A*B Total Sales Revenue $1,292,013
D Input Price per unit $110.00
E=A*D Total Input cost -$444,840
F Unit Variable Cost $73.48
G=A*F Total Variable Costs -$297,163
H Fixed costs -$638,978
I Depreciation -$207,668
J=C+E+G+H+I Before tax operating income -$296,636
K=J*30% Tax expenses $88,991
L=J+K After tax Operating Income -$207,645
M Add:Depreciation (non cash expense) $207,668
N=L+M Operating Cash Flow $23
Break even sales in units for operating cash flow 4044
c Operating Cash Flow for sales of 8000 units
Operating Cash Flow(Dollar) Dollar
A Number of units sold 8000
B Sales Price per unit $319.49
C=A*B Total Sales Revenue $2,555,911
D Input Price per unit $110.00
E=A*D Total Input cost -$880,000
F Unit Variable Cost $73.48
G=A*F Total Variable Costs -$587,859
H Fixed costs -$638,978
I Depreciation -$207,668
J=C+E+G+H+I Before tax operating income $241,406
K=J*30% Tax expenses -$72,422
L=J+K After tax Operating Income $168,984
M Add:Depreciation (non cash expense) $207,668
N=L+M Operating Cash Flow $376,652
d If spot rate reduces to RMB 4.6/USD
Operating Cash Flow(Dollars) Dollar
A Number of units sold 7000
B Sales Price per unit $434.78 (2000/4.6)
C=A*B Total Sales Revenue $3,043,478
D Input Price per unit $110.00
E=A*D Total Input cost -$770,000
F Unit Variable Cost $100.00 (460/4.6)
G=A*F Total Variable Costs -$700,000
H Fixed costs -$869,565 (4million/4.6)
I Depreciation -$282,609 (1.3 million/4.6)
J=C+E+G+H+I Before tax operating income $421,304
K=J*30% Tax expenses -$126,391
L=J+K After tax Operating Income $294,913
M Add:Depreciation (non cash expense) $282,609
N=L+M Operating Cash Flow $577,522
New Operating Cash Flow=$577,522
Reduction in Sales Price to obtain Operating Cash Flow= $281,447
Operating Cash Flow(Dollars) Dollar
A Number of units sold 7000
B Sales Price per unit $374.36 (2000/4.6)
C=A*B Total Sales Revenue $2,620,520
D Input Price per unit $110.00
E=A*D Total Input cost -$770,000
F Unit Variable Cost $100.00 (460/4.6)
G=A*F Total Variable Costs -$700,000
H Fixed costs -$869,565 (4million/4.6)
I Depreciation -$282,609 (1.3 million/4.6)
J=C+E+G+H+I Before tax operating income -$1,654
K=J*30% Tax expenses $496
L=J+K After tax Operating Income -$1,158
M Add:Depreciation (non cash expense) $282,609
N=L+M Operating Cash Flow $281,451
New Selling Price in RMB =374.36*4.6= $1,722.06 RMB
Reduction in price =2000-1722.06 $277.94 RMB